# Science and Technology

### v1

@@ 5.4 Ablation Study
To determine the effectiveness of our proposed method, we first conduct ablation studies under the image-goal setting of the InstanceNav task. In Table 4, we report the evaluation results of different variants of our method. We start from the ZSON [37] baseline, with a fixed CLIP goal encoder and a learnable ResNet-50 observation encoder. Then we apply the entropy-minimized goal view selection algorithm to select meaningful goal images during pre-training data pre-processing. This modification brings $+2.1\%$ improvement in success rate. We then add the perspective reward relaxation technique, which enhances goal orientation with an alternative pitch angle. This approach aims to prevent the agent from being immersed in exactly matching the goal view and observations, neglecting the semantic correlation between them. However, we discovered that with the vanilla model design, the augmented image goal confuses the agent and decreases the navigation success rate (from 12.7% to 10.8%). We argue that this phenomenon is due to the inferior agent design that neglects to reason semantic differences in the goal images and observation images. Incorporating with the PSL agent alleviates this issue. We found that it boosts the navigation success rate by $+3.8\%$ with the goal view selection approach, but more impressively, it cooperates well with the perspective relaxation mechanism and contributes a lot to the improvement ($+9.3\%$ in SR and $+4.2\%$ in SPL). We attribute it to the semantic perception module. This module captures and compares semantics in both goal and observation images with the same granularity, preventing the agent from learning sub-optimal navigation policies.
Method Input PSL Agent PST Strategy SR SPL
SPM GVS PRR
ZSON [37] Image 12.7 6.5
PSL (Ours) Image 14.8 7.7
Image 10.8 5.5
Image 16.5 6.5
Image 22.0 10.7
@@
5.5 Qualitative Analysis
Qualitative examples for PSL agent.
In Fig. 6 we present qualitative examples of our PSL agent navigating to two different object instances given detailed language instruction (e.g., , “Find the plant made of the black pot and grayish-white branches, around with a transparent glass dining table and four white chairs.”). The agent navigates across rooms, determining relative object instances according to the instruction, and finally stops near the object. We find that given detailed language instruction, the agent is able to differentiate between objects of the same category. For instance, in the first case, the agent bypasses a plant with “a white bottle and pink flower” which differs from the intrinsic attribute description. After a left turn near the first plant, the agent arrives at its destination and faces the correct plant instance with “black pot and grayish-white branches”. In another case, the agent hovers for a while in the intersection of bedroom and dining room to find the “white chair” and last the episode at “several white chairs with a glass dining table” that mentioned in the extrinsic attributes.
Figure 6: Qualitative examples for our PSL agent navigating to an object instance according to intrinsic and extrinsic object attributes in the InstanceNav dataset. For each trail, the agent is initialized at a random position in the room and given language instruction “Find the …”.
Figure 7: t-SNE vis. of the embeddings used by semantic expansion inference scheme.
Visualization of semantic expansion inference.
In Fig. 7, we provide t-SNE visualization of the embeddings used in our semantic expansion inference scheme. We present a triplet example where we retrieve an embedding from the support set using a text query. The retrieved embedding shows less difference from the ground-truth goal image embedding compared to the original text query. Our support set encompasses the object categories in InstanceNav and extends to a broader range of embeddings from various perspectives, contributing to the performance gain in navigation success rate.
@@ 6 Conclusion
We present PSL, a zero-shot approach for instance navigation. PSL is built upon the open-vocabulary and cross-modal ability of CLIP, and targets addressing the limited semantic perceiving ability of the observation encoder in existing methods. Our method is composed of three different parts, including a PSL agent with a semantic perception module module to reason the semantic differences between the observation and goal, a prioritized semantic training strategy with clear semantic supervision to train the PSL agent, and a semantic expansion inference scheme to maintain the granularity level of the goal-semantic. We investigate how this agent acts when instructed to navigate to a specific object by extending the zero-shot object navigation task to a zero-shot instance navigation task in the widely-used HM3D environment. we find that most existing methods struggle in this task, while our PSL agent outperforms other approaches.

### v2

@@ 5.4 Ablation Study
We first conduct ablation studies under the image-goal setting of the InstanceNav task. We start from the ZSON [38] baseline, with a fixed CLIP goal encoder and a learnable ResNet-50 observation encoder. Then we apply the entropy-minimized goal view selection algorithm to select meaningful goal images during pre-training data pre-processing. In Table 4, this modification brings $+2.1\%$ improvement in success rate. We then add the perspective reward relaxation technique, which enhances goal orientation with an alternative pitch angle. This approach aims to prevent the agent from being immersed in exactly matching the goal view and observations, neglecting the semantic correlation between them. However, we discovered that with the vanilla model design, the augmented image goal confuses the agent and decreases the navigation success rate (from 12.7% to 10.8%). We argue that this phenomenon is due to the inferior agent design that neglects to reason semantic differences in the goal images and observation images. Incorporating with the PSL agent alleviates this issue. We found that it boosts the navigation success rate by $+3.8\%$ with the goal view selection approach, but more impressively, it cooperates well with the perspective relaxation mechanism and contributes a lot to the improvement ($+9.3\%$ in SR and $+4.2\%$ in SPL).
@@
PSL PST ZSIN-image ZSIN-text ZSON
SPM GVS PRR SR SPL SR SPL SR SPL
ZSON 12.7 6.5 10.6 6.5 25.5 12.6
PSL (Ours) 19.5 7.9 13.0 5.6 33.7 15.8
14.8 7.7 11.8 6.1 30.4 14.7
16.5 6.5 12.3 5.7 35.0 18.1
22.0 10.7 16.5 7.5 42.4 19.2
@@ 6 Conclusion
We present PSL, a zero-shot approach for instance navigation. PSL is built upon the open-vocabulary and cross-modal ability of CLIP, and targets addressing the limited semantic perceiving ability of existing methods. Our method is composed of three different parts, including a PSL agent with a semantic perception module module to reason the semantic differences between the observation and goal, a prioritized semantic training strategy with clear semantic supervision to train the PSL agent, and a semantic expansion inference scheme to maintain the granularity of the goal-semantic. We investigate how this agent acts when instructed to navigate to a specific object by extending the zero-shot object navigation task to a zero-shot instance navigation task. We find that most existing methods struggle in this task, while our PSL agent outperforms other approaches.
Acknowledgements
This work was supported by the Meituan Academy of Robotics Shenzhen, the Guangzhou Municipal Science and Technology Project (No. 2024A03J0619) and the National Natural Science Foundation of China (No. 62306257).